Have to write the equation for question 3 and 4
Adrian Dingles Chemistry Pages and ChemEducator LLC 2013 All rights rese 3 The standard enthalpy of combustion of H%) The standard enthalpy of combustion of (s) Hesss Law Hesss Law states that, the enthalpy change during a depends only on the na reaction reactants and products and is independent of the route taken. Task 5Civ Using one of the following methods to calculate the enthalpy changes questions
Answer #1

Q1.

STD enthalpy for combustion of H2(G)

so

H2(g) + O2(g) = H2O(g)

balance

H2(g) + 1/2O2(g) = H2O(g)

Hcomb = Hprod - Hreact = H2O - (H2 + 1/2O2)

Q4.

Combusiton of Al(s)

Al(s) + O2(g) = Al2O3(s)

balance

2Al(s) + 3/2O2(g) = Al2O3(s)

Hcomb = Al2O3 - (2*Al + 3/2*O2)